OpenStudy (wayward_son):

Write the equation of a line that passes through the points (0, -2) and (4, -5). Answer must be in standard form: Ax + By = C. Please show work. ☺

OpenStudy (mathmale):

First of all, you need to find the slope of the straight line connecting the 2 given points. Know the slope formula? Secondly, choose either the slope-intercept form y=mx+b or the point-slope form\[y-y _{?}=m(x-x _{1})\]

OpenStudy (mathmale):

In either case, substitute your numerical value for m into the equation (replacing m). Now choose either one of the 2 given points. This point has an x-component and a y-component. Subst. these 2 numerical values into your equation (whichever one you chose), above.

OpenStudy (mathmale):

WaywardSon: You seem wayward. ;) Looking around?

OpenStudy (mathmale):

Sorry. I have other things to do right now, if you're busy with something else. Work with you another time. Good luck.

OpenStudy (mathmale):

You might get more help and more satisfaction from OpenStudy if you'd stick around for at least a few minutes after having posted a problem. Welcome to OpenStudy!

OpenStudy (marihelenh):

If you still need help with this problem, let me know and I will work through it with you.

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)